OpenStudy (anonymous):

help please--- 2(x-2)=4x-2

OpenStudy (anonymous):

do u understand, please help me

Nnesha (nnesha):

first you have to distribute x -2 by 2

OpenStudy (anonymous):

yes

Nnesha (nnesha):

so what you get

OpenStudy (anonymous):

2x-4

Nnesha (nnesha):

ok so now add 4x both side

OpenStudy (anonymous):

2x-4x=4x-2

Nnesha (nnesha):

nope

OpenStudy (anonymous):

2x-4x+4x=-2+4+4x

Nnesha (nnesha):

after distributing you get this

OpenStudy (anonymous):

yea

Nnesha (nnesha):

there is now 4x on left side

Nnesha (nnesha):

no**

Nnesha (nnesha):

now subtract 4x both side